# Masked Attribute Description Embedding for Cloth-Changing Person Re-identification

### v2

@@ Abstract
Cloth-changing person re-identification (CC-ReID) aims to match persons who change clothes over long periods. The key challenge in CC-ReID is to extract clothing-independent features, such as face, hairstyle, body shape, and gait. Current research mainly focuses on modeling body shape using multi-modal biological features (such as silhouettes and sketches). However, it does not fully leverage the personal description information hidden in the original RGB image. Considering that there are certain attribute descriptions which remain unchanged after the changing of cloth, we propose a Masked Attribute Description Embedding (MADE) method that unifies personal visual appearance and attribute description for CC-ReID. Specifically, handling variable clothing-sensitive information, such as color and type, is challenging for effective modeling. To address this, we mask the clothing and color information in the personal attribute description extracted through an attribute detection model. The masked attribute description is then connected and embedded into Transformer blocks at various levels, fusing it with the low-level to high-level features of the image. This approach compels the model to discard clothing information. Experiments are conducted on several CC-ReID benchmarks, including PRCC, LTCC, Celeb-reID-light, and LaST. Results demonstrate that MADE effectively utilizes attribute description, enhancing cloth-changing person re-identification performance, and compares favorably with state-of-the-art methods. The code is available at https://github.com/moon-wh/MADE.
1 Introduction
@@
Person re-identification (Re-ID) refers to the process of identifying persons across different camera views. Traditional Re-ID methods operate under the assumption that the person being tracked moves within a confined area and time and will not change their clothes. However, in practical scenarios, persons captured by surveillance cameras may traverse larger areas and be observed over extended periods, during which they might change their clothing. This deviation in appearance challenges the reliability of color-based information utilized by earlier Re-ID approaches for person re-identification. Therefore, in recent years, cloth-changing person re-identification has attracted more and more attention.
@@
In cloth-changing person re-identification, the reliance on color information by traditional methods becomes unreliable, and addressing the modeling of clothing changes poses a significant challenge [44]. Therefore, the key to solving CC-ReID lies in identifying information about personal clothing that is insensitive to these changes. To mitigate the interference caused by varying clothes and uncover invariant features of persons, some cloth-changing person re-identification methods focus on the multi-modal biometric features of persons. PRCC [44] and FSAM [19] study silhouette information of persons. 3DSL [4] aims to learn the 3D shape features of persons. However, contours and 3D features eliminate all color information of images. MBUNet [47] contains a branch for extracting posture features. GI-ReID [23] and ViT-VIBE [2] utilize gait features. However, extracting pose features from a single image of a person is still a challenging task. SpTSkM [48] uses skeleton normalization to assist person recognition. RF-ReID [11] infers personal skeleton features from radio frequency signals. However, skeleton features are challenging to extract and utilize directly. These methods often require additional complex models for extracting multi-modal biometric features, demanding substantial computing resources for training. This complexity overlooks the most direct description information present in personal images. Moreover, the way these methods utilize the extracted clothing-insensitive information is usually more complex. CAL [15] and the method proposed by chan et al. [3] are both based on GAN [14] network.
Figure 1: In the scene of cloth-changing person re-identification, the person captured across cameras has many attributes that remain unchanged (attribute words marked in blue) except for their clothes (attribute words marked in black).
@@
In the context of cloth-changing person re-identification, even over extended time intervals, persons tend to alter only their clothing choices, while other attributes such as gender, age, and hair color remain consistent, as depicted in Fig. 1. These clothes-interrelated attributes are helpful to person re-identification, while clothes-related attributes can be easily eliminated. This paper introduces the Masked Attribute Description Embedding (MADE) method to effectively mine clothing-insensitive information from original RGB images of persons. Specifically, we adopt the attribute detection model to extract pedestrian attributes. Then, the clothing and color information is masked to obtain masked attribute description, and the clothing-sensitive features are eliminated by shielding the clothing information in the RGB image and retaining other clothing-insensitive color features. Due to the editable nature of text description, clothing color can be quickly and efficiently eliminated. MADE then connects and embeds the masked attributes description data encoded by Linear Projection into TrV blocks [12] at different levels to fuse the image features. By mapping different feature spaces to a shared latent space, masked description can be fused with image features without the need for additional text encoders, forcing the model to discard clothing-sensitive information.
@@
  • We introduce multi-modal attribute description information in CC-ReID that is more evident and easier to extract and edit than skeletons or contours in the original person’s images. Since the clothing-sensitive information is not conducive to identifying persons, we use the attribute detection model to extract the description in Description Extraction and Mask(DEM) module, mask the variable clothing and color items and embed it into the image features, which can force the model to discard clothing-sensitive features.
  • For the first time, we employ a simple and efficient method to integrate image features with attribute description in CC-ReID. MADE connects and embeds the masked attribute description data, encoded by Linear Projection, into transformer blocks at various levels, facilitating the fusion of low-level and high-level image features. Through the mapping of different feature spaces to a shared latent space, the model gains a deeper understanding and exploitation of semantic relationships. Without the need for additional text encoders, descriptions seamlessly fuse with image features, enabling the model to capture the associated information between images and descriptions effectively.
@@
In the following, we will discuss related work in section 2. We will present the details of our poposed method in section 3. The experimental results are provided in section 4. Section 5 concludes this paper with our future reseach directions
2 Related Work
2.1 Cloth-Changing Person Re-Identification
The core of solving cloth-changing person re-identification is to extract the clothes-irrelevant features in person images. To this end, some research focuses on multi-modal biometric features that are less variable than clothing, such as silhouette, 3D shape, skeleton, walking posture, etc.
@@ 2.1 Cloth-Changing Person Re-Identification
SpTSkM [48] explores persons’ motion pattern information from 3D skeletons normalized by ST-GAN [43] to assist person re-identification. 3DSL [4] distinguishes different identities by learning 3D shape features and 3D reconstruction subnetworks. These methods obtain 3D shapes through cumbersome 3D parsing and processing networks, which increases the complexity of model training.
@@
CESD [31] uses a pose detector to detect personal body joint points and uses shape embedding to separate clothing and distinguish shape information through joint point features. The pose feature branch of MBUNet  [47] applies the direction adaptive graph convolution layer to obtain the relevant information between different keypoints in heatmaps. ViT-VIBE [2] uses ViT [10] to combine appearance and gait features learned through VIBE [25]. However, these methods do not fully exploit and utilize the clothing-independent features in the original image.
Some methods use adversarial strategies to mine clothing-insensitive information in original images. AFD-Net [42] produces more realistic images by using a discriminator for data augmentation. CCFA [16] uses an adversarial loss to supervise the augmentation vector to ensure that it roughly satisfies the empirical clothing change distribution in the training data. CAL [15] classifies clothing based on a GAN network and penalizes the model’s ability to predict clothing. However, adversarial models are prone to training instability and have difficulty explaining how to separate appearance and clothing information from RGB images.
@@
In this paper, we propose MADE to unify persons’ appearance and language description. Multi-modal attribute description information is introduced in CC-ReID, which is more obvious and accessible to extract and edit than biological features such as skeleton or silhouette in original person images. Information that helps identify persons can be retained to the greatest extent while accurately removing interference from clothing information.
2.2 Text-to-Image Person Retrieval
@@
Li et al[27] created the first benchmark dataset CUHK-PEDES. The main challenge in this field is how to efficiently fuse image and text features into a joint embedding space. Early research work [26, 27] adopted VGG [38] and LSTM [18] to learn the representation of visual-text modalities. Subsequent studies [7, 35, 49] designed new cross-modal matching losses to align global image-text features in the joint embedding space. IRRA [22] integrates visual cues with CLIP [33] encoded text tokens into a cross-modal multi-modal interaction encoder, enabling cross-modal interaction. These methods utilize sentence captions describing persons. However, when dealing with scenes involving changes in clothing, the model faces challenges in directly processing clothing-related fragments within the captions. Introducing an encoder to encode language text would increase the complexity of model training. To address this, we leverage itemized attributes. This approach enables the model to precisely handle clothing-related information within the description, avoiding the processing of the entire statement as a whole. Simultaneously, we embed the attribute vector directly into the Transformer block, eliminating the need for additional encoder encoding.
2.3 Attribute based Person Re-identification
@@
It has been well exploited to perfrom person re-identification with attributes. APR [28] introduces the Attribute Reweighting Module (ARM), which corrects predictions of attributes based on learned dependencies and correlations between attributes. MSPA [24] uses ConvLSTM to memorize the relationship between persons’ attribute features. These methods utilize all pedestrian attributes to address traditional person re-identification challenges but overlook the editability of the description. We introduce the editability of description in CC-ReID, which can accurately remove clothing-sensitive attributes and help the model learn clothing-sensitive information.
3 Method
3.1 Overview
@@
The key to achieving cloth-changing person re-identification lies in extracting clothing-insensitive features from the image. In CC-ReID, we introduce attribute description to mitigate the impact of clothing interference. Consequently, our model primarily focuses on modeling the relationship between and within the two types of images and descriptions. EVA-02 [12] is pre-trained to reconstruct powerful and robust language-aligned visual features through occlusion image modeling, resulting in transferable models. Based on this model, we proposed MADE framework to integrate persons’ masked attribute description data with image visual features, addressing the challenges of cloth-changing person re-identification. The framework of our approach is shown in Fig. 2. Given an image sample $x_{i}$, $x_{i}$ extracts an editable attribute description through Description Extraction and Mask module. After the masked attribute description is converted into a binary vector, it is connected and embedded at different levels through Linear Projection to fuse with image features in TrV blocks [12]. We introduce how to extract and mask attribute description in Sec. 3.2. Then, Sec. 3.3 details how to add masked attribute description data to improve the performance of CC-ReID. Finally, we elaborate on the model’s loss function and inference process.
@@
3.2 Description Extraction and Mask (DEM)
In CC-ReID, it is imperative for the model to disregard clothing information in the RGB image during input to learn clothing-insensitive features. DeSKPro [41] and SAVS [13] use the human parsing model to generate person parsing maps to remove clothing interference. However, processing the parsing map to obtain robust cloth-irrelevant information is more complicated. The description information of persons mainly describes the appearance and clothing of persons, so previous research rarely involves processing person description in CC-ReID. However, we could eliminate interference information from person’s clothes in the model input by simply editing the attribute description. In MADE, for each input image $x_{i}$, we extract the description and operate the clothes mask through DEM, as shown in Fig. 2.
DEM uses the extraction model to obtain persons’ descriptions suitable for cloth-changing datasets and performing mask processing. Specifically, we use SOLIDER [6] trained on PETA [9] to identify person attributes in cloth-changing datasets. SOLIDER [6] is a human task visual pre-training model that adopts self-supervised training. We use it to obtain attribute description of images. PETA contains 19,000 images, including 8,705 individuals, each annotated with 61 binary and four multiclass attributes, 105 attribute labels in total. Given a sample $x_{i}$, input SOLIDER to get persons’ attributes list, [$a_{1}^{i}$, $a_{2}^{i}$,…,$a_{N}^{i}$], and convert them into 0-1 binary vectors by position. We implement the clothing information mask operation by setting the clothing attributes to 0. And, the masked attribute description data [$m_{1}^{i}$, $m_{2}^{i}$,…,$m_{N}^{i}$] with the clothing information removed is obtained.
@@ 3.2 Description Extraction and Mask (DEM)
3.3 Process of Description Embedding
@@
Given an image sample $x_{i}\in R^{H\times W\times C}$and its masked-description data $m_{i}$ = [$m_{1}^{i}$, $m_{2}^{i}$,…,$m_{N}^{i}$], we integrate them into TrV Block [12] to get the MADE framework, as shown in Figure 2. TrV Block is an improved Vision Transformer structure. We use EVA02-large as the backbone, which has 24 layers of TrV Blocks and divides it into three stages (we discuss it in detail in 4.6.2). First, we segment the image $x_{i}$ into a sequence of $N=H\times W/P^{2}$ fixed-size, where P represents the size of the patch, and then map the patch sequence through a trainable linear projection as one-dimensional notation {${f_{j}^{v}|_{j=1}^{N}}$}. After the injection of positional embedding and additional [CLS] token, the tokens sequence {$f_{cls}^{v},f_{1}^{v},…,f_{N}^{v}$} is input into the TrV block of $L$ layers of the first stage to model dependencies between each patch. Subsequently, $f_{cls}^{v}$ is extracted to represent the first stage’s global image low-level feature representation. Then $m_{i}$ is passed through Linear Projection to obtain the description feature {${f_{j}^{m}|_{j=1}^{N}}$} with the same dimension as {${f_{j}^{v}|_{j=1}^{N}}$}. After adding the extra token [DES], the attribute description sequence is {$f_{des}^{m},f_{1}^{m},…,f_{N}^{m}$}. The description sequence is then connected to the image sequence, {$f_{j}^{vm}|_{j=1}^{N}$}={$f_{des}^{m},f_{1}^{m},…,f_{N}^{m},f_{1}^{v},…,f_{N}^{v}$}, input the second and third stages for training. In this process, image features and description features are embedded through connections and trained together to learn the relationship between images and attribute description that mask clothing information. The interference of clothing-sensitive can be removed through mask items and connection embedding, avoiding the problem of complex extraction and fusion of multi-modal biometric features. In the model, we extract $f_{des}^{m}$ as a fusion representation of image and attribute description features. The class tokens output by the second and third stages, $f_{des,2}^{m}$ and $f_{des,3}^{m}$, respectively, represent the fusion of different levels of visual features and attribute description of clothing removal. We combine them with the output of $f_{cls}^{v}$ from the first stage to obtain the pedestrian feature representation of MADE through Conv1d aggregation.
3.4 Loss Function and Inference

### v3

@@ Abstract
Cloth-changing person re-identification (CC-ReID) aims to match persons who change clothes over long periods. The key challenge in CC-ReID is to extract cloth-irrelated features, such as face, hairstyle, body shape, and gait. Current research mainly focuses on modeling body shape using multi-modal biological features (such as silhouettes and sketches). However, it does not fully leverage the personal description information hidden in the original RGB image. Considering that there are certain attribute descriptions that remain unchanged after the changing of cloth, we propose a Masked Attribute Description Embedding (MADE) method that unifies personal visual appearance and attribute description for CC-ReID. Specifically, handling variable cloth-sensitive information, such as color and type, is challenging for effective modeling. To address this, we mask the clothes type and color information (upper body type, upper body color, lower body type, and lower body color) in the personal attribute description extracted through an attribute detection model. The masked attribute description is then connected and embedded into Transformer blocks at various levels, fusing it with the low-level to high-level features of the image. This approach compels the model to discard cloth information. Experiments are conducted on several CC-ReID benchmarks, including PRCC, LTCC, Celeb-reID-light, and LaST. Results demonstrate that MADE effectively utilizes attribute description, enhancing cloth-changing person re-identification performance, and compares favorably with state-of-the-art methods. The code is available at https://github.com/moon-wh/MADE.
I Introduction
@@
Cloth-changing person re-identification (CC-ReID) aims to match persons who change clothes over long periods. Traditional person re-identification (Re-ID) operates under the assumption that the person being tracked moves within a confined area and time and will not change their clothes. However, in practical scenarios, persons captured by surveillance cameras may traverse larger areas and be observed over extended periods, during which they might change their clothes. This deviation in appearance challenges the reliability of color-based information utilized by earlier Re-ID approaches for person re-identification. Therefore, in recent years, cloth-changing person re-identification has attracted more and more attention.
@@
In cloth-changing person re-identification, the reliance on color information by traditional methods becomes unreliable, and addressing the modeling of clothes changes poses a significant challenge [1]. Therefore, the key to solving CC-ReID lies in identifying information about personal features that is insensitive to these clothes changes. To mitigate the interference caused by varying clothes and uncover invariant features of persons, some cloth-changing person re-identification methods focus on the multi-modal biometric features of persons. PRCC [1] and FSAM [2] study silhouette information of persons. 3DSL [3] aims to learn the 3D shape features of persons. However, contours and 3D features eliminate all color information of images. MBUNet [4] contains a branch for extracting posture features. GI-ReID [5] and ViT-VIBE [6] utilize gait features. However, extracting pose features from a single image of a person is still a challenging task. SpTSkM [7] uses skeleton normalization to assist person recognition. RF-ReID [8] infers personal skeleton features from radio frequency signals. However, skeleton features are challenging to extract and utilize directly. CAL [9] and the method proposed by Chan et al. [10] do not utilize multi-modal biometric features; instead, they use GAN [11] networks to extract cloth-irrelated features from pedestrian images. However, GAN networks have unstable training, long training times, and difficulty tuning hyperparameters. These methods typically require additional complex models to extract biometric features such as contours, 3D shapes, postures, and skeletons, which demand substantial computing resources for training and extraction and also require complex fusion of the extracted biometric features with image features. Methods that do not use multi-modal features often face issues with unstable training.
Figure 1: An illustration of cloth-changing person re-identification over a long period of time and across cameras. The attributes of the person image is shown in the figure. Attributes related to clothes are marked in black, while attributes irrelevant to clothes are marked in blue. In the cloth-changing person re-identification scenario, many attributes unrelated to clothes remain consistent, such as hair, glasses, shoes, age, and gender, which could be useful for re-identification.
@@
In the context of cloth-changing person re-identification, even over extended time intervals, persons tend to alter only their clothes choices, while other attributes such as gender, age, and hair color remain consistent, as depicted in Fig. 1. This cloth-irrelevant attributes are helpful to person re-identification, while cloth-related attributes can be easily eliminated. This paper introduces the Masked Attribute Description Embedding (MADE) method to effectively mine cloth-irrelevant information from original RGB images of persons. Specifically, we adopt the attribute detection model SOLIDER [12] to extract pedestrian attributes. SOLIDER is a self-supervised learning framework used to learn universal human representations from a large number of unannotated human images. It demonstrates excellent performance on pedestrian attribute recognition tasks. Then, by masking the cloth-related pedestrian attributes to obtain masked attribute descriptions (the definition of cloth-related attributes is in section III-B), and the cloth-sensitive features are eliminated by shielding the cloth information in the RGB image and retaining other cloth-insensitive color features. Due to the editable nature of text descriptions, clothes color can be quickly and efficiently eliminated. MADE then connects and embeds the masked attributes description data encoded by Linear Projection into TrV blocks [13] at different levels to fuse the image features. By mapping different feature spaces to a shared latent space, masked description can be fused with image features without the need for additional text encoders, forcing the model to discard cloth-sensitive information.
@@
  2. 2.
    We introduce multi-modal attribute description information in CC-ReID, which is easier to extract and edit than skeletons or contours. By masking clothes and cloth-color items in these descriptions and embedding them into image features, the model is compelled to discard cloth-sensitive features.
  3. 3.
    For the first time, we employ a simple, efficient method to integrate image features with attribute descriptions in CC-ReID. This approach maps descriptions and image features to a shared latent space, effectively allowing the model to capture their associations without additional text encoders.
  4. 4.
@@
In the following, we will discuss related work in section II. We will present the details of our proposed method in section III. The experimental results are provided in section IV. Section V concludes this paper with our future research directions
II Related Work
II-A Multi-Modal Features based Cloth-Changing Person Re-Identification
The core of solving cloth-changing person re-identification is to extract the cloth-irrelevant features in person images. To this end, some research focuses on multi-modal features that are less variable than clothes, such as silhouette, 3D shape, skeleton, walking posture, etc.
@@ II-A Multi-Modal Features based Cloth-Changing Person Re-Identification
SpTSkM [7] explores personal motion pattern information from 3D skeletons normalized by ST-GAN [16] to assist person re-identification. 3DSL [3] distinguishes different identities by learning 3D shape features and 3D reconstruction subnetworks. These methods obtain 3D shapes through cumbersome 3D parsing and processing networks, which increases the complexity of model training.
@@
CESD [17] uses a pose detector to detect personal body joint points and uses shape embedding to separate clothes and distinguish shape information through joint point features. The pose feature branch of MBUNet  [4] applies the direction adaptive graph convolution layer to obtain the relevant information between different keypoints in heatmaps. ViT-VIBE [6] uses ViT [18] to combine appearance and gait features learned through VIBE [19]. However, these methods do not fully exploit and utilize the cloth-irrelevant features in the original image.
@@
In this paper, we propose MADE to unify personal appearance and language description. Multi-modal attribute description information is introduced in CC-ReID, which is more obvious and accessible to extract and edit than biological features such as skeleton or silhouette in original person images. Information that helps identify persons can be retained to the greatest extent while accurately removing interference from cloth information.
II-B Text-to-Image Person Re-Identification
@@
Text-to-image person re-identification aims to search for pedestrian images of an interested identity via textual descriptions. The main challenge in this field is how to efficiently fuse image and text features into a joint embedding space. Early research work [20, 21] adopted VGG [22] and LSTM [23] to learn the representation of visual-text modalities. CFine [24] proposes a CLIP-driven fine-grained information excavation framework to fully utilize the powerful knowledge of CLIP for text-image person re-identification. IRRA [25] integrates visual cues with CLIP [26] encoded text tokens into a cross-modal multi-modal interaction encoder, enabling cross-modal interaction. These methods utilize sentence captions describing persons. However, when dealing with scenes involving changes in clothes, the model faces challenges in directly processing cloth-related fragments within the captions. Introducing an encoder to encode language text would increase the complexity of model training. To address this, we leverage itemized attributes. This approach enables the model to precisely handle cloth-related information within the description, avoiding the processing of the entire statement as a whole. Simultaneously, we embed the attribute vector directly into the Transformer block, eliminating the need for additional encoder encoding.
II-C Attribute-based Person Re-identification
@@
It has been well exploited to perform person re-identification with attributes. APR [27] introduces the Attribute Reweighting Module (ARM), which corrects predictions of attributes based on learned dependencies and correlations between attributes. AAB [28] utilizes fine-grained attribute attention modules to enhance the performance of the Re-ID task. MSPA [29] uses ConvLSTM to memorize the relationship between personal attribute features. AMNet [30] designs the Spatial Channel Attention Module (SCAM) to extract features from each attribute. Additionally, it utilizes the semantic reasoning and information propagation capabilities of graph convolutional networks to explore the relationship between attribute features and pedestrian features. UCAD [31] proposes a clothes attribute decomposition network that can effectively attenuate the influence of clothes through loss function constraints. These methods utilize all pedestrian attributes to address person re-identification challenges but overlook the editability of the attributes description. We introduce the editability of description in CC-ReID, which can accurately remove cloth-sensitive attributes and help the model learn cloth-sensitive information.
III Method
III-A Overview
@@
The key to achieving cloth-changing person re-identification lies in extracting cloth-insensitive features from the image. In CC-ReID, we introduce attribute description to mitigate the impact of clothes interference. Consequently, our model primarily focuses on modeling the relationship between and within the two types of images and descriptions. EVA-02 [13] is pre-trained to reconstruct powerful and robust language-aligned visual features through occlusion image modeling, resulting in transferable models. Based on this model, we proposed MADE framework to integrate personal masked attribute description data with image visual features, addressing the challenges of cloth-changing person re-identification. The framework of our approach is shown in Fig. 2. Given an image sample $x_{i}$, $x_{i}$ extracts an editable attribute description through Description Extraction and Mask module. After the masked attribute description is converted into a binary vector, it is connected and embedded at different levels through Linear Projection to fuse with image features in TrV blocks [13]. We introduce how to extract and mask attribute description in section III-B. Then, section III-C details how to add masked attribute description data to improve the performance of CC-ReID. Finally, we elaborate on the model’s loss function and inference process.
@@
III-B Description Extraction and Mask (DEM)
In CC-ReID, it is imperative for the model to disregard cloth information in the RGB image during input to learn cloth-insensitive features. DeSKPro [32] and SAVS [33] use the human parsing model to generate person parsing maps to remove clothes interference. However, processing the parsing map to obtain robust cloth-irrelevant information is more complicated. The description information of persons mainly describes the appearance and clothes of persons, so previous research rarely involves processing person description in CC-ReID. However, we could eliminate interference information from person’s clothes in the model input by simply editing the attribute description. In MADE, for each input image $x_{i}$, we extract the description and operate the clothes mask through DEM, as shown in Fig. 2.
DEM uses the extraction model to obtain personal descriptions suitable for cloth-changing datasets and performing mask processing. Specifically, we use SOLIDER [12] trained on PETA_ZS [34] to identify person attributes in cloth-changing datasets. SOLIDER [12] is a human task visual pre-training model that adopts self-supervised training. We use it to obtain attribute descriptions of images. PETA_ZS contains 19,000 images, including 8,705 individuals, each annotated with 61 binary and four multiclass attributes, 105 attribute labels in total. They can be divided into categories such as gender, age, orientation, type of carried items, upper body color, upper body type, lower body color, lower body type, shoe color, and shoe type. We define upper body color, upper body type, lower body color, and lower body type as cloth-related attributes, while the others are defined as cloth-unrelated attributes. Given a sample $x_{i}$, input SOLIDER to get personal attributes list, [$a_{1}^{i}$, $a_{2}^{i}$,…,$a_{N}^{i}$], and convert them into 0-1 binary vectors by position. We implement the cloth information mask operation by setting the clothes attributes to 0. And, the masked attribute description data [$m_{1}^{i}$, $m_{2}^{i}$,…,$m_{N}^{i}$] with the cloth information removed is obtained.
Fig. 3 presents some examples of pedestrian attributes extracted using SOLIDER [12]. Attribute recognition is a multi-classification task, where the attributes recognized for each pedestrian are not precisely the same. Among these attributes, those related to clothes (upper body color, upper body type, lower body color, and lower body type) are marked in black, while attributes unrelated to clothes are marked in blue. The cloth-related attributes are masked (set to 0) to obtain the final masked description data.
@@ III-B Description Extraction and Mask (DEM)
Figure 3: Examples of pedestrian attribute lists extracted using SOLIDER (Attributes related to clothes are marked in black, while attributes unrelated to clothes are marked in blue).
III-C Description Embedding
@@
Given an image sample $x_{i}\in R^{H\times W\times C}$and its masked-description data $m_{i}$ = [$m_{1}^{i}$, $m_{2}^{i}$,…,$m_{N}^{i}$], we integrate them into TrV Block [13] to get the MADE framework, as shown in Fig. 2. Considering the accuracy of the attribute extraction model, we introduce random 0-1 noise into the masked-description data (the discussion regarding the correctness of the attribute extraction model and the proportion of noise added is in section IV-E1). TrV Block is an improved Vision Transformer structure. We use EVA02-large as the backbone, which has 24 layers of TrV Blocks and divides it into three stages (we discuss it in detail in section IV-E3). First, we segment the image $x_{i}$ into a sequence of $N=H\times W/P^{2}$ fixed-size, where P represents the size of the patch, and then map the patch sequence through a trainable linear projection as one-dimensional notation {${f_{j}^{v}|_{j=1}^{N}}$}. After the injection of positional embedding and additional [CLS] token, the tokens sequence {$f_{cls}^{v},f_{1}^{v},…,f_{N}^{v}$} is input into the TrV block of $L$ layers of the first stage to model dependencies between each patch. Subsequently, $f_{cls}^{v}$ is extracted to represent the first stage’s global image low-level feature representation.
Then masked-description data $m_{i}$ is passed through Linear Projection and expanded to three dimensions to obtain the description feature {$f^{m}$} aligning the third dimension with the same dimension as {$f_{j}^{v}\mid_{j=1}^{N}$}. After adding the extra token [DES], the attribute description sequence is {$f_{des}^{m},f^{m}$}, as shown in Fig. 2. The description sequence is then connected to the image sequence, {$f_{j}^{vm}\mid_{j=1}^{N}$} = {$f_{des}^{m},f^{m},f_{1}^{v},\ldots,f_{N}^{v}$}, and input into the second and third stages for training. In this process, image features and description features are embedded through connections and trained together to learn the relationship between images and attribute descriptions that mask cloth information. The interference of cloth-sensitive features can be removed through mask items and connection embedding, avoiding the problem of complex extraction and fusion of multi-modal biometric features.
In the model, we extract $f_{des}^{m}$ as a fusion representation of image and attribute description features. The class tokens output by the second and third stages, $f_{des,2}^{m}$ and $f_{des,3}^{m}$, respectively, represent the fusion of different levels of visual features and attribute description of clothes removal. We combine them with the output of $f_{cls}^{v}$ from the first stage to obtain the pedestrian feature representation of MADE through Conv1d aggregation.
III-D Loss Function and Inference